# On computability of equilibrium states

### v2

@@ 4.2. Proof of the computability of the Jacobian
In this subsection, we aim to establish uniform computability of the Jacobian, as stated in the following theorem.
@@
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, on the following input
@@
    two algorithms computing the potential $\varphi$ and the map $T$, respectively,
@@
    an oracle of a point $x\in X$,
@@
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$,
@@
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
@@
    a constant $n\in\mathbb{N}$,
the algorithm outputs a rational $2^{-n}$-approximation for the value of $J_{\varphi}(x)$, where the function $J_{\varphi}$ is the function defined by
(4.19) $J_{\varphi}(x)\coloneqq\frac{u_{\varphi}(T(x))}{u_{\varphi}(x)}\exp(P(T,\varphi)-\varphi(x)),\qquad x\in X.$
@@
Consider $x\in X$ and an oracle $\varphi$ for $x$. By Definition 3.8, $\{x\}=\bigcap_{n\in\mathbb{N}}\overline{B(s_{\varphi(n)},2^{-n})}$. Hence $T^{-1}(x)=\bigcap_{n\in\mathbb{N}}T^{-1}\bigl{(}\overline{B(s_{\varphi(n)},2^{-n})}\bigr{)}$. By Definition 3.6, we get that $\bigl{\{}\bigl{(}\overline{B(s_{\varphi(n)},2^{-n})}\bigr{)}^{c}\bigr{\}}_{n\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets. Since $T$ is computable, it follows from Proposition 3.10 that the sequence $\bigl{\{}T^{-1}\bigl{(}\bigl{(}\overline{B(s_{\varphi(n)},2^{-n})}\bigr{)}^{c}\bigr{)}\bigr{\}}_{n\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets. By Proposition 3.7, $\bigl{(}T^{-1}(x)\bigr{)}^{c}=\bigcup_{n\in\mathbb{N}}T^{-1}\bigl{(}\bigl{(}\overline{B(s_{\varphi(n)},2^{-n})}\bigr{)}^{c}\bigr{)}$ is a lower-computable open set. By Proposition 3.15, noting that $X$ is recursively compact, $T^{-1}(x)$ is a recursively compact set. By [GHR10, Proposition 4], $X$ is recursively precompact, and consequently, there is an algorithm that outputs an $\eta$-net of $X$. Because $T$ is distance-expanding with the constants $\eta$, $\lambda$, and $\xi$, there is at most one preimage point of $x$ in each ball with radius $\eta$. Therefore each preimage point of $x$ is computable. ∎
@@
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the value of $\mathcal{L}^{m}_{\varphi}(\mathbbm{1})(x)$, after inputting the following data in this algorithm:
  1. (i)
    two algorithms computing the potential $\varphi$ and the transformation $T$, respectively,
  2. (ii)
  3. (iii)
    an oracle of a point $x\in X$,
  4. (iv)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  5. (v)
    two constants $n,\,m\in\mathbb{N}$.
We can now apply Theorem 4.1 to prove the computability of the pressure $P(T,\varphi)$ and the function $u_{\varphi}(x)$ (see Corollary 2.9 for the definition of the function $u_{\varphi}(x)$).
@@
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the topological pressure $P(T,\varphi)$, after inputting the following data in this algorithm:
  1. (i)
    two algorithms computing the potential $\varphi$ and the map $T$, respectively,
  2. (ii)
  3. (iii)
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$,
  4. (iv)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  5. (v)
    a constant $n\in\mathbb{N}$.
@@
    Compute a $\xi^{\prime}$-net $G$ for the space $X$, where $\xi^{\prime}\coloneqq\min\{\eta,\,\xi\}$. Then compute a constant $N\in\mathbb{N}$ satisfying that $\bigcap_{x\in G}\bigcup_{k=0}^{N}T^{k}(x,\xi)=X.$
@@
    Find $N_{1}\in\mathbb{N}$ with $2^{n+1}N\log(\operatorname{card}G)\cdot\bigl{(}\frac{4a_{0}\xi^{v_{0}}}{\lambda^{v_{0}}-1}+2N\|\varphi\|_{\infty}\bigr{)}\leqslant N_{1}$.
@@
    $v_{n}\approx w_{n}\coloneqq N_{1}^{-1}\log\bigl{(}\mathcal{L}_{\varphi}^{N_{1}}(\mathbbm{1})(x_{0})\bigr{)}$
    with precision $2^{-n-1}$, where $x_{0}\in\mathcal{S}$ is an ideal point.
Let us verify that $v_{n}$ satisfies $\lvert v_{n}-P(T,\varphi)\rvert<2^{-n}$ for each $n\in\mathbb{N}$. To see this, it suffices to check that $\lvert w_{n}-P(T,\varphi)\rvert<2^{-n-1}$ for each $n\in\mathbb{N}$. Since $T\colon X\rightarrow X$ is distance-expanding with constants $\eta$, $\lambda,$ and $\xi$, then $T$ is injective on each balls of radius $\xi^{\prime}_{0}$. Hence, we have $D\coloneqq\max_{x\in X}\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}\leqslant\operatorname{card}G$. Moreover, by Lemma 2.8, and steps (4) and (5) above, one can conclude that
$\lvert w_{n}-P(T,\varphi)\rvert=\bigl{\lvert}N_{1}^{-1}\log\bigl{(}e^{-{N_{1}}P(T,\varphi)}\mathcal{L}_{\varphi}^{N_{1}}(\mathbbm{1})(x_{0})\bigr{)}\bigr{\rvert}<(\log C)/N_{1}<2^{-n-1}.\qed$
@@
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the value of $u_{\varphi}(x)$, after inputting the following data in this algorithm:
  1. (i)
    two algorithms computing the potential $\varphi$ and the map $T$, respectively,
  2. (ii)
  3. (iii)
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$,
  4. (iv)
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
  5. (v)
    an oracle of a point $x\in X$,
  6. (vi)
    a constant $n\in\mathbb{N}$.
@@
    Compute a $\xi^{\prime}$-net $G$ for the space $X$, where $\xi^{\prime}\coloneqq\min\{\eta,\,\xi\}$. Then compute a constant $N\in\mathbb{N}$ satisfying that $\bigcap_{x\in G}\bigcup_{k=0}^{N}T^{k}(x,\xi)=X.$
@@
[…]
    $\overline{C}\coloneqq(\operatorname{card}G)^{N}\exp\Bigl{(}\frac{4a_{0}\xi^{v_{0}}}{\lambda^{v_{0}}-1}+2N\|\varphi\|_{\infty}\Bigr{)}.$
@@
    Use Theorem 4.1 to compute $\epsilon\in(0,\xi)$ with $2\exp(a^{\prime}\epsilon^{v_{0}})<C^{\prime}.$ Then compute a $(\epsilon/2)$-net $G^{\prime}$ for the space $X$, and a constant $m\in\mathbb{N}$ satisfying that $\bigcap_{t\in G^{\prime}}T^{m}(t,\epsilon/2)=X$.
@@
    Find $k\in\mathbb{N}$ satisfying that
@@
    $s_{n}\approx t_{n}\coloneqq e^{-mkP(T,\varphi)}\mathcal{L}_{\varphi}^{mk}(\mathbbm{1})(x)$
@@
[…]
$\lvert t_{n}-u_{\varphi}(x)\rvert\leqslant\bigl{\|}\mathcal{L}_{\overline{\varphi}}^{mk}(\mathbbm{1})(x)-u_{\varphi}(x)\bigr{\|}_{\infty}\leqslant CeZ\bigl{(}1-e^{-Z}\bigr{)}^{k-1}<\overline{C}e\overline{Z}\bigl{(}1-e^{-\overline{Z}}\bigr{)}^{k-1}<2^{-n-1},$
@@ 4.3. Two subsets of $\mathcal{P}(X)$
In this subsection, we consider two subsets of $\mathcal{P}(X)$ closely related to the equilibrium state $\mu_{\varphi}$ and prove that they are both lower-computable open sets.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $X$ is recursively compact, and $T\colon X\rightarrow X$ is computable. Then the set $U$ of Borel probability measures that are not $T$-invariant is a lower-computable open set.
@@
Let $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be a computable enumeration of $\mathcal{E}(\mathcal{S})$ (see Definition 3.22) in the computable metric space $(X,\,\rho,\,\mathcal{S})$. Claim. $U=\bigcup_{j\in\mathbb{N}}\Phi_{j}$, where $\Phi_{j}\coloneqq\left\{\mu\in\mathcal{P}(X):\int\!\varphi_{j}\,\mathrm{d}\mu-\int\!\varphi_{j}\circ T\,\mathrm{d}\mu\neq 0\right\}$ for each $j\in\mathbb{N}$. Indeed, consider $\mu\in\bigcap_{j\in\mathbb{N}}\Phi_{j}^{c}$. By Proposition 3.23, we conclude that $\mu=\mu\,\circ\,T^{-1}$, namely, $\mu\notin U$. Hence, $U=\bigcup_{j\in\mathbb{N}}\Phi_{j}$, establishing the claim. Next, note that $\{\varphi_{j}\}_{j\in\mathbb{N}}$ is a uniformly computable sequence of functions and $T$ is computable. Then by Corollary 3.20, the family $\{\mathcal{I}_{j}\}_{j\in\mathbb{N}}$ of integral operators $\mathcal{I}_{j}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ defined by
$\mathcal{I}_{j}(\mu)\coloneqq\int\!(\varphi_{j}-\varphi_{j}\circ T)\,\mathrm{d}\mu$
is a uniformly computable sequence of functions. By Definition 3.5, $\mathbb{R}\setminus\{0\}$ is a lower-computable open set. Note that $\Phi_{j}=\mathcal{I}_{j}^{-1}(\mathbb{R}\setminus\{0\})$. Then by Corollary 3.11, $\{\Phi_{j}\}_{j\in\mathbb{N}}$ is a uniformly computable sequence of lower-computable open sets. Therefore, by Proposition 3.7 and the claim, $U$ is a lower-computable open set. ∎
@@
Let the quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfy the Assumption A in Section 1. Define the set $W$ so that a Borel probability measure $\mu$ is in $W$ if and only if the Jacobian of $T\colon X\rightarrow X$ with respect to $\mu$ is not
@@
Let $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be a computable enumeration of $\mathcal{E}(\mathcal{S})$ (see Definition 3.22) in the computable metric space $(X,\,\rho,\,\mathcal{S})$, and $\{x_{k}:1\leqslant k\leqslant m\}$ be an $\eta$-net for $X$. Set $g_{k,n}(x)\coloneqq g_{x_{k},\eta-\frac{1}{n},\frac{1}{n}}(x)$ (see Definition 3.21) for each $n\in\mathbb{N}$ and each integer $k\leqslant m$. Claim. $W=\bigcup_{k=1}^{m}\bigcup_{j,\,n\in\mathbb{N}}\Phi_{j}^{k,n}$, where
@@
for all $j,\,k,\,n\in\mathbb{N}$ with $k\leqslant m$. By the definition of $W$ and Jacobians, $\Phi_{j}^{k,n}\subseteq W$ for all $j,\,k,\,n\in\mathbb{N}$ with $k\leqslant m$. Therefore, to prove the claim, it suffices to show $W\subseteq\bigcup_{k=1}^{m}\bigcup_{j,\,n\in\mathbb{N}}\Phi_{j}^{k,n}$. Fix an integer $1\leqslant k\leqslant m$. Since $T$ is distance-expanding with constants $\eta$, $\lambda,$ and $\xi$, it is injective on $B(x_{k},\eta)$. Thus, the function $(\varphi_{j}\cdot g_{k,n})\circ T^{-1}$ is well-defined in $X$ for all $j,\,n\in\mathbb{N}$. Consider
@@
Then for each $j\in\mathbb{N}$, by (4.21), we have
$\int\!(\varphi_{j}\cdot g_{k,n})\circ T^{-1}\,\mathrm{d}\mu=\int\!\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}\,\mathrm{d}\mu\quad\text{for each }n\in\mathbb{N}.$
Letting $n\to+\infty$, it follows from Dominated Covergence Theorem that
@@
for each $j\in\mathbb{N}$. Then by Proposition 3.23, it follows that $\mu_{1}=\mu_{2}$, where $\mu_{1}$ and $\mu_{2}$ are given by
@@
for each Borel set $A\subseteq X$. Thus,
$\mu(T(A\cap B(x_{k},\eta)))=\int_{A\cap B(x_{k},\eta)}\!J_{\varphi}\,\mathrm{d}\mu$
for each Borel set $A\subseteq X$. Therefore, for each $\mu\in\bigcap\limits_{1\leqslant k\leqslant m}\bigcap\limits_{j,\,n\in\mathbb{N}}\bigl{(}\Phi^{k,n}_{j}\bigr{)}^{c}$ and Borel set $A\subseteq X$, we can define $A_{1}\coloneqq A\cap B(x_{1},\eta)$, and $A_{k}\coloneqq\bigl{(}A-\bigcup_{i=1}^{k-1}A_{i}\bigr{)}\cap B(x_{k},\eta)$ for each integer $k\in[2,m]$. Since $X=\bigcup_{k=1}^{m}B(x_{k},\eta)$, then $A=\bigsqcup_{k=1}^{m}A_{k}$. Hence,

### v3

@@ 4.2. Proof of the computability of the Jacobian
In this subsection, we establish the uniform computability of Jacobians, as stated in the following theorem.
@@
There exists an algorithm with the following property:
For each $x_{0}\in X$, each $n\in\mathbb{N}$, and each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, the algorithm outputs a rational $2^{-n}$-approximation for the value of $J_{\varphi}(x_{0})$, where the function $J_{\varphi}\colon X\rightarrow\mathbb{R}$ is defined by
(4.20) $J_{\varphi}(x)\coloneqq\frac{u_{\varphi}(T(x))}{u_{\varphi}(x)}\exp(P(T,\varphi)-\varphi(x))\quad\text{ for each }x\in X,$
on the following input:
@@
    two algorithms computing $\varphi$ and $T$, respectively,
@@
    two rational constants $a_{0}>0$ and $v_{0}\in(0,1]$ with $\varphi\in C^{0,v_{0}}(X)$ and $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$,
@@
    three rational constants $\eta>0$, $\lambda>1$, and $\xi>0$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
@@
    an oracle of the point $x_{0}\in X$,
@@
    the constant $n\in\mathbb{N}$.
@@
For each $p\in X$, denote by $\overline{B}(p,r)\coloneqq\{q\in X:\rho(p,q)\leqslant r\}$ the closed ball of radius $r$. Consider a point $x\in X$. By [GHR11, Proposition 4], we obtain that $X$ is recursively precompact, and consequently, there is an algorithm outputting an $\eta$-net of $X$, say $\{x_{i}:i\in[1,n]\cap\mathbb{N}\}$. Because $T$ is distance-expanding with constants $\eta$, $\lambda$, and $\xi$, by (2.3), there is at most one preimage point of $x$ in each closed ball with radius $\eta$. By Definition 3.14, $\overline{B}(x_{i},\eta)$ is recursively compact for each integer $1\leqslant i\leqslant n$. Hence, by Proposition 3.15 (iii), the set $B_{i}=\overline{B}(x_{i},\eta)\smallsetminus\bigcup_{j=1}^{i-1}B(x_{j},\eta)$ is recursively compact for each integer $1\leqslant i\leqslant n$. Thus by [BRY12, Theorem 2.12], the inverse $\bigl{(}T|_{B_{i}}\bigr{)}^{-1}\colon T(B_{i})\rightarrow B_{i}$ of the restriction of $T$ is computable for each $i\in\mathbb{N}$. Therefore, the inverse $T^{-1}$ of $T$ is computable (in the sense of Definition 3.13). ∎
@@
There exists an algorithm with the following property:
For each $x_{0}\in X$, each pair of $n,\,m\in\mathbb{N}$, and each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the value of $\mathcal{L}^{m}_{\varphi}(\mathbbm{1})(x_{0})$, after inputting an oracle of the point $x_{0}$, the constants $n,\,m\in\mathbb{N}$, and data (i) through (iv) from Theorem 4.7 in this algorithm.
Now we apply Lemma 2.8 to establish the computability of the value of the pressure $P(T,\varphi)$.
@@
There exists an algorithm with the following property:
For each $n\in\mathbb{N}$ and each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the topological pressure $P(T,\varphi)$, after inputting the constant $n\in\mathbb{N}$ and data (i) through (iv) from Theorem 4.7 in this algorithm.
@@
    Compute a $\xi^{\prime}$-net $G$ for the space $X$, where $\xi^{\prime}\coloneqq\min\{\eta,\,\xi\}$. Then compute $N\in\mathbb{N}$ with $\bigcap_{x\in G}\bigcup_{k=0}^{N}T^{k}(x,\xi)=X$.
@@
    Compute $N_{1}\in\mathbb{N}$ with $2^{n+1}N\log(\operatorname{card}G)\cdot\bigl{(}\frac{4a_{0}\xi^{v_{0}}}{\lambda^{v_{0}}-1}+2N\|\varphi\|_{\infty}\bigr{)}\leqslant N_{1}$.
@@
    $v_{n}\approx w_{n}\coloneqq N_{1}^{-1}\log\bigl{(}\mathcal{L}_{\varphi}^{N_{1}}(\mathbbm{1})(y_{0})\bigr{)}$
    with precision $2^{-n-1}$, where $y_{0}\in\mathcal{S}$ is an ideal point.
Let us verify that $v_{n}$ satisfies $\lvert v_{n}-P(T,\varphi)\rvert<2^{-n}$ for each $n\in\mathbb{N}$. To see this, it suffices to check that $\lvert w_{n}-P(T,\varphi)\rvert<2^{-n-1}$ for each $n\in\mathbb{N}$. Since $T\colon X\rightarrow X$ is distance-expanding with constants $\eta$, $\lambda,$ and $\xi$, then $T$ is injective on each balls of radius $\xi^{\prime}$. Hence, we have $D\coloneqq\max_{x\in X}\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}\leqslant\operatorname{card}G$. Moreover, by Lemma 2.8, and steps (4) and (5) above, we can conclude that
$\lvert w_{n}-P(T,\varphi)\rvert=\bigl{\lvert}N_{1}^{-1}\log\bigl{(}e^{-{N_{1}}P(T,\varphi)}\mathcal{L}_{\varphi}^{N_{1}}(\mathbbm{1})(y_{0})\bigr{)}\bigr{\rvert}<(\log C)/N_{1}<2^{-n-1}.\qed$
Applying Theorem 4.1, we can show that the function $u_{\varphi}\colon X\rightarrow\mathbb{R}$ is computable.
@@
There exists an algorithm with the following property:
For each $x_{0}\in X$, each $n\in\mathbb{N}$, and each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a rational $2^{-n}$-approximation for the value of $u_{\varphi}(x_{0})$, after inputting an oracle of the point $x_{0}$, the constant $n\in\mathbb{N}$, and data (i) through (iv) from Theorem 4.7 in this algorithm.
@@
    Compute a $\xi^{\prime}$-net $G$ for the space $X$, where $\xi^{\prime}\coloneqq\min\{\eta,\,\xi\}$. Then compute a constant $N\in\mathbb{N}$ satisfying $\bigcap_{x\in G}\bigcup_{k=0}^{N}T^{k}(x,\xi)=X.$
@@
    Use Theorem 4.1 to compute a constant $C^{\prime}\in\mathbb{R}$ satisfying $C^{\prime}>\max\bigl{\{}\overline{C}^{2},\,2\exp(a^{\prime})\bigr{\}}$, where
    (4.21) $\overline{C}\coloneqq(\operatorname{card}G)^{N}\exp\Bigl{(}\frac{4a_{0}\xi^{v_{0}}}{\lambda^{v_{0}}-1}+2N\|\varphi\|_{\infty}\Bigr{)}.$
@@
    Use Theorem 4.1 to compute $\epsilon\in(0,\xi)$ with $2\exp(a^{\prime}\epsilon^{v_{0}})<C^{\prime}.$ Then compute an $(\epsilon/2)$-net $G^{\prime}$ for the space $X$, and a constant $m\in\mathbb{N}$ satisfying $\bigcap_{t\in G^{\prime}}T^{m}(t,\epsilon/2)=X$.
@@
    Find $k\in\mathbb{N}$ satisfying
@@
    $s_{n}\approx t_{n}\coloneqq e^{-mkP(T,\varphi)}\mathcal{L}_{\varphi}^{mk}(\mathbbm{1})(x_{0})$
@@
Fix an integer $n$. Now we verify that $s_{n}$ satisfies $\lvert s_{n}-u_{\varphi}(x_{0})\rvert<2^{-n}$. To see this, it suffices to check that $\lvert t_{n}-u_{\varphi}(x_{0})\rvert<2^{-n-1}$. Since $T\colon X\rightarrow X$ is distance-expanding with constants $\eta$, $\lambda,$ and $\xi$, $T$ is injective on each ball of radius $\xi^{\prime}$. Hence, we have that $D\coloneqq\max_{x\in X}\operatorname{card}\bigl{(}T^{-1}(x)\bigr{)}\leqslant\operatorname{card}G$. Recall that in Lemma 2.8 we define
$C\coloneqq(\operatorname{card}G)^{N}\exp\biggl{(}\frac{4a_{0}\xi^{v_{0}}}{\lambda^{v_{0}}-1}+2N\|\varphi\|_{\infty}\biggr{)}.$
So by (4.21), $\overline{C}\geqslant C$. Thus $C^{\prime}>\max\bigl{\{}\overline{C}^{2},\,2\exp(a^{\prime})\bigr{\}}\geqslant\max\bigl{\{}C^{2},\,2\exp(a^{\prime})\bigr{\}}$. Moreover, by (4.1) and Step (6) above, we obtain that
$\overline{Z}\geqslant Z\coloneqq\widehat{K}\biggl{(}\frac{\lambda^{-v_{0}}+1}{2},C^{\prime},\frac{(C^{-2}Q^{m}+2)C^{\prime}}{2C^{-2}Q^{m}+2}\biggr{)}.$
By Theorem 4.1, it follows from $0<Z(1-\exp(-Z))^{k-1}\leqslant\overline{Z}(1-\exp(-\overline{Z}))^{k-1}\leqslant 1$ that
$\lvert t_{n}-u_{\varphi}(x_{0})\rvert\leqslant\bigl{\|}\mathcal{L}_{\overline{\varphi}}^{mk}(\mathbbm{1})(x_{0})-u_{\varphi}(x_{0})\bigr{\|}_{\infty}\leqslant CeZ\bigl{(}1-e^{-Z}\bigr{)}^{k-1}\leqslant\overline{C}e\overline{Z}\bigl{(}1-e^{-\overline{Z}}\bigr{)}^{k-1}<2^{-n-1},$
@@ 4.3. Two subsets of $\mathcal{P}(X)$
In this subsection, we consider two subsets of $\mathcal{P}(X)$ which are closely related to the equilibrium state $\mu_{\varphi}$ and prove that they are both lower-computable open sets.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space. Assume that $X$ is recursively compact, and $T\colon X\rightarrow X$ is computable. Then $U\coloneqq\mathcal{P}(X)-\mathcal{M}(X,T)$ is a lower-computable open set.
@@
Define the function $L_{T}\colon\mathcal{P}(X)\rightarrow\mathcal{P}(X)$ by duality in the following way: for each $\mu\in\mathcal{P}(X),\,L_{T}(\mu)$ satisfies that $\int\!f\,\mathrm{d}L_{T}(\mu)=\int\!(f\circ T)\,\mathrm{d}\mu$ for each $f\in C(X)$. Since the map $T\colon X\rightarrow X$ is computable, by [GHR11, Theorem 3.1], the map $L_{T}\colon\mathcal{P}(X)\rightarrow\mathcal{P}(X)$ is also computable. Define $\mathcal{I}\colon\mathcal{P}(X)\rightarrow\mathbb{R}$ by $\mathcal{I}(\mu)=W_{\rho}(\mu,\,L_{T}(\mu))$ for each $\mu\in\mathcal{P}(X)$. Then $\mathcal{I}$ is a computable function. By Definition 3.6, $\mathbb{R}\smallsetminus\{0\}$ is a lower-computable open set. Hence, by Proposition 3.10, $\mathcal{I}^{-1}(\mathbb{R}\smallsetminus\{0\})$ is a lower-computable open set. By the definition of $U$, $U=\mathcal{I}^{-1}(\mathbb{R}\smallsetminus\{0\})$. Therefore, $U$ is a lower-computable open set. ∎
@@
Let the quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfy the Assumption A in Section 1. Define the set $W$ so that a Borel probability measure $\mu$ is in $W$ if and only if $J_{\varphi}$ is not the Jacobian of $T\colon X\rightarrow X$ with respect to $\mu$, where
@@
Let $\{\varphi_{j}\}_{j\in\mathbb{N}}$ be a computable enumeration of $\mathfrak{E}(\mathcal{S})$ (see Definition 3.22) in the computable metric space $(X,\,\rho,\,\mathcal{S})$, and $\{x_{k}:k\in[1,m]\cap\mathbb{N}\}$ be an $\eta$-net for $X$. Set $g_{k,n}(x)\coloneqq g_{x_{k},\eta-\frac{1}{n},\frac{1}{n}}(x)$ (see Definition 3.21) for all $k,\,n\in\mathbb{N}$ with $k\leqslant m$.
Claim. $W=\bigcup_{k=1}^{m}\bigcup_{j,\,n\in\mathbb{N}}\Phi_{j}^{k,n}$, where
@@
for all $j,\,k,\,n\in\mathbb{N}$ with $k\leqslant m$.
By the definitions of $W$ and the function $J_{\varphi}$, we have $\Phi_{j}^{k,n}\subseteq W$ for all $j,\,k,\,n\in\mathbb{N}$ with $k\leqslant m$. Hence, to prove the claim, it suffices to show that $W\subseteq\bigcup_{k=1}^{m}\bigcup_{j,\,n\in\mathbb{N}}\Phi_{j}^{k,n}$. Fix an integer $1\leqslant k\leqslant m$. Since $T$ is distance-expanding with constants $\eta$, $\lambda,$ and $\xi$, it is injective on $B(x_{k},\eta)$. Thus, the function $(\varphi_{j}\cdot g_{k,n})\circ T^{-1}$ is well-defined in $X$ for all $j,\,n\in\mathbb{N}$.
Consider
@@
Then for each $j\in\mathbb{N}$ and each $n\in\mathbb{N}$, by (4.23), we have
$\int\!(\varphi_{j}\cdot g_{k,n})\circ T^{-1}\,\mathrm{d}\mu=\int\!\varphi_{j}\cdot g_{k,n}\cdot J_{\varphi}\,\mathrm{d}\mu.$
Fix $j\in\mathbb{N}$ and let $n\to+\infty$. It follows from the Dominated Convergence Theorem that
@@
for each $j\in\mathbb{N}$. Then by Proposition 3.23, we have $\mu_{1}=\mu_{2}$, where $\mu_{1}$ and $\mu_{2}$ are given by
@@
for each Borel set $A\subseteq X$. Thus, $\mu(T(A\cap B(x_{k},\eta)))=\int_{A\cap B(x_{k},\eta)}\!J_{\varphi}\,\mathrm{d}\mu$ for each Borel set $A\subseteq X$.
Hence, for each $\mu\in\bigcap\limits_{1\leqslant k\leqslant m}\bigcap\limits_{j,\,n\in\mathbb{N}}\bigl{(}\Phi^{k,n}_{j}\bigr{)}^{c}$ and each Borel set $A\subseteq X$, we can define $A_{1}\coloneqq A\cap B(x_{1},\eta)$, and $A_{k}\coloneqq\bigl{(}A-\bigcup_{i=1}^{k-1}A_{i}\bigr{)}\cap B(x_{k},\eta)$ for each integer $k\in[2,m]$. Since $X=\bigcup_{k=1}^{m}B(x_{k},\eta)$, then $A=\bigsqcup_{k=1}^{m}A_{k}$. So